# Nondeterministic Polynomial-time Problem Challenge: An Ever-Scaling Reasoning Benchmark for LLMs

### v1

@@ Abstract
Reasoning is the fundamental capability of large language models (LLMs). Due to the rapid progress of LLMs, there are two main issues of current benchmarks: i) these benchmarks can be crushed in a short time (less than 1 year), and ii) these benchmarks may be easily hacked. To handle these issues, we propose the ever-scalingness for building the benchmarks which are uncrushable, unhackable, auto-verifiable and general. This paper presents Nondeterministic Polynomial-time Problem Challenge (NPPC), an ever-scaling reasoning benchmark for LLMs. Specifically, the NPPC has three main modules: i) npgym, which provides a unified interface of 25 well-known NP-complete problems and can generate any number of instances with any levels of complexities, ii) npsolver: which provides a unified interface to evaluate the problem instances with both online and offline models via APIs and local deployments, respectively, and iii) npeval: which provides the comprehensive and ready-to-use tools to analyze the performances of LLMs over different problems, the number of tokens, the aha moments, the reasoning errors and the solution errors. Extensive experiments over widely-used LLMs demonstrate: i) NPPC can successfully decrease the performances of advanced LLMs’ performances to below 10%, demonstrating that NPPC is uncrushable, ii) DeepSeek-R1, Claude-3.7-Sonnet, and o1/o3-mini are the most powerful LLMs, where DeepSeek-R1 can outperform Claude-3.7-Sonnet and o1/o3-mini in most NP-complete problems considered, and iii) the numbers of tokens, aha moments in the advanced LLMs, e.g., Claude-3.7-Sonnet and DeepSeek-R1, are observed first to increase and then decrease when the problem instances become more and more difficult. We believe that NPPC is the first ever-scaling reasoning benchmark of LLMs, serving as the uncrushable and unhackable testbed for LLMs toward artificial general intelligence (AGI).
@@ 1 Introduction
The remarkable successes of Large Language Models (LLMs) [1] have catalyzed the fundamental shift of artificial intelligence. The recent breakthrough on reasoning [14] enables the LLMs to complete complex tasks, e.g., math proof, code generation and computer use, which require the capabilities of understanding, generation and long-term planning. Various benchmarks, e.g., GPQA [24], AIME, SWE-bench [17] and ARC-AGI [6], are proposed to evaluate these advanced reasoning capabilities, where most benchmarks are curated and verified by human researchers with a finite number of questions. However, there are two main issues of current benchmarks: i) current benchmarks can be crushed in a short time (as shown in Figure 1): the performance on GSM8K [8] is increased from about 35% to 95% in three years, while the performance on SWE-bench [17] is increased from 7.0% to 64.6% in about 8 months, and ii) current benchmarks can be easily hacked or exploited: ChatbotArena leverages human votes to evaluate the LLMs, which may cost about \$ 3000 for one evaluation, while the MixEval [21] can reduce the cost to less than \$ 1 with high correlation. This low-cost evaluation brings the possibility of hacking ChatbotArena. The two main issues bring difficulties for the evaluation of LLMs and emerge as the main obstacles in a rapidly progressing era of LLMs.
Figure 2: Ever-scalingness
To build a successful benchmark for LLMs, we propose the ever-scalingness with the following four desiderata for a benchmark: i) uncrushable, which requires the scaling over the complexity, i.e., the benchmark can generate the problems with continually increasing complexities, ii) unhackable, which requires the scaling over instances, i.e., the benchmark can generate an infinite number of instances to avoid the exploitation, iii) auto-verifiable, which requires the scaling over the scalable oversight, i.e., the benchmark can verify the correctness of the solutions efficiently for the problems with any complexity, and iv) general, which requires the scaling over the coverage, i.e., the problems covered by the benchmark should be highly relevant to the real-world problems, rather than some puzzles or rare problems.
Figure 3: Motivation of NPPC
To build the ever-scaling benchmark, we consider nondeterminstic polynomial-time (NP) problems, i.e., the problems where the solutions can be verified in polynomial time [10]. Specifically, we focus on the NP-complete (NPC) problems, which are the hardest ones among all NP problems. The main reasons for choosing NPC problems for our ever-scaling benchmark are: i) the instances of NPC problems can be systematically generated with any level of difficulty, e.g., combinatorial search space, thus leading to the scaling of complexity and instances, ii) NPC problems are intrinsic to be “difficult to solve, easy to verify” and we do not have any polynomial-time algorithms to solve NPC problems. Therefore, the NPC problems are uncrushable even by any algorithms or tools so far and the solutions of any instance of NPC problems can be verified efficiently, i.e., scaling of the scalable oversight, and iii) many real-world problems, e.g., navigation and planning, can be formulated as NP(C) problems and any other NP problems can be transformed (or, reduced) to NPC problems in polynomial time. NPC problems are the foundation problems of all computational problems and LLMs are the foundation models for wide range tasks, thus leading to the emergence of our ever-scaling nondeterministic polynomial-time problem challenge (NPPC) (as displayed in Figure 3).
@@
Specifically, the NPPC has three main modules: i) npgym, which provides a unified interface of 25 well-known NPC problems and can generate any number of instances with any levels of complexities, which implies the ever-scalingness of NPPC, ii) npsolver, which provides a unified interface to evaluate the problem instances with both online and offline models via APIs and local deployments, respectively, to facilitate users to evaluate their own models and iii) npeval, which provides comprehensive and ready-to-use tools to analyze the performances of LLMs over different problems, the number of tokens, the “aha moments”, the reasoning errors and the solution errors, which can provide in-depth analysis of the LLMs and the insights to further improve the LLMs’ reasoning capabilities. Extensive experiments over widely-used LLMs, i.e., GPT-4o-mini, GPT-4o, Claude-3.7-Sonnet, DeepSeek-V3, DeepSeek-R1, and OpenAI o1-mini, demonstrate: i) NPPC can successfully decrease the performances of advanced LLMs’ performances to below 10%, demonstrating that NPPC is uncrushable, ii) DeepSeek-R1, Claude-3.7-Sonnet, and o1/o3-mini are the most powerful LLMs, where DeepSeek-R1 can outperform Claude-3.7-Sonnet and o1-mini in most NP-complete problems considered, and iii) the numbers of tokens, aha moments in the advanced LLMs, e.g.. Claude-3.7-Sonnet and DeepSeek-R1, are observed to first increase and then decrease when the problem instances become more and more difficult. We also analyze the different errors of reasoning and the solutions in the LLMs and the issues of offline models for solving complex reasoning problems. To the best of our knowledge, NPPC is the first ever-scaling reasoning benchmark of LLMs, serving as the uncrushable and unhackable testbed for the advanced LLMs toward artificial general intelligence (AGI). The code is released at https://github.com/SMU-DIGA/nppc.
@@ 2 Related Work
Uncrush.
Unhack.
Auto- verify
General
NPHardEval [13]
ZebraLogic [20]
Reasoning Gym [22]
Sudoku-Bench [25]
ARC-AGI-1 & 2 [6]
@@
We will provide a review of existing reasoning benchmarks in this section. Abstraction and Reasoning Corpus (ARC-AGI)-1 [6] is designed to be “easy for humans, hard for AI”, which is formed by human-curated 800 puzzle-like tasks, designed as grid-based visual reasoning problems. ARC-AGI-1 are featured by OpenAI as the leading benchmark to measure the performance of their o3 models. o3 at low compute scored 75.7% on ARC-AGI-1 and reached 87% accuracy with higher compute, which roughly crushes the ARC-AGI-1 benchmarks and leads to the emergence of the ARC-AGI-2 benchmark. ARC-AGI-1 & 2 are the representative of traditional benchmarks for LLMs, e.g., MMLU-Pro [28], GPQA [24], GSM8K [8], and SWE-bench [17], which are formed by static or regularly updated questions curated and verified by human. We believe that these benchmarks will never be uncrushable and may be hacked by some techniques.
Several recent benchmarks consider either NP(C) problems, e.g., 3SAT [4, 15, 23], or partially the ever-scalingness [13, 22], (displayed in Table 1). NPHardEval [13] considers 3 problems from P, NPC and NP-hard classes and use these class to evaluate the LLMs. We note that the problems in P class can be solved by augmenting the LLMs with tools, e.g., code running, and the NP-hard problems cannot be verified efficiently, therefore, NPHardEval cannot scale over the scalable oversight. Only 3 NPC problems are considered, i.e., Knapsack, Traveling salesman problem (TSP) and graph coloring, and the instances of each problem in NPHardEval are finite and only regularly updated, which cannot scale over the instance and complexity. ZebraLogic [20] considers one logic puzzle, i.e., Zebra puzzle, to test the reasoning capabilities of LMs when the problems’ complexities increase. However, the reasoning capability on specific puzzles does not necessarily transfer to other problems, which violates the scaling of the coverage. Sudoku-Bench [25] focuses on one specific Sudoku game with 2765 procedurally generated instances with various difficulty levels. Reasoning Gym [22] is an ongoing project which collects the procedural generators and algorithmic verifiers for infinite training data with adjustable complexity. Though with some NP(C) problems, e.g., Zebra puzzles and Sudoku, the reasoning gym does not specifically focus on NPC problems and scaling over complexity.
@@ 3.1 P, NP and NP-complete Problems
P and NP Problems.
The problems in P class are decision problems that can be solved in polynomial time by a deterministic Turing machine, which implies there exists an algorithm that can find a solution in time proportional to a polynomial function, e.g., $O(n^{k})$, of the input size $n$. Examples include sorting, shortest path problems, and determining if a number is prime. The problems in NP class are decision problems that can be solved in polynomial time by nondeterministic Turing machine, where a proposed solution can be easily verified, though finding that solution might require more time (as displayed in Definition 1). All P problems are also in NP, but the reverse remains an open question, known as “P vs. NP problem”. NP problems form the cornerstone of computational complexity theory, for which solution verification is tractable (polynomial time) even though solution discovery may be intractable (potentially exponential time), i.e., “difficult to solve, easy to verify”. Many real-world optimization problems can be formulated as NP problems, such as equilibrium finding in game theory, portfolio management, network design and machine learning.
@@
NP-complete (NPC) Problems.
Formally, a problem $\Omega$ is an NPC problem if i) the problem is in NP, and ii) any NP problems can be transformed to problem $\Omega$ in polynomial time. This reducibility property establishes NPC problems as the "hardest" problems in NP class. The Cook-Levin theorem established SAT as the first proven NPC problem [9, 18], while 3SAT is the special case of SAT and is also an NPC problem. Subsequent NPC problems typically proven via reduction chains back to 3SAT or other established NPC problems. The most well-known NPC problems include vertex cover problem, clique problem, traveling salesman proble (TSP), Hamiltonian path/cycle problem, etc. NPC problems play the most important roles in answering the “P vs. NP problem”, i.e., if any NPC problem were shown to have a polynomial-time algorithm, then P = NP. However, despite decades of research, no polynomial-time algorithms for any NPC problems have been discovered, which implies that the NPC problems are uncrushable by current methods or algorithms.
3.2 Reasoning in LLMs
The reasoning ability of LLMs refers to the model’s capacity to process information in a systematic way, which enables LLMs to tackle complex problems, e.g., mathematical proof and code generation, that require multi-step thinking, context understanding, and knowledge integration. Recently, specialized reasoning models have been proposed. OpenAI-o1 is an LLM trained with reinforcement learning (RL), which enables the model to perform complex reasoning, including logical thinking and problem solving, via chain-of-thought (CoT). o1 thinks before it answers and can significantly outperform GPT-4o on reasoning-heavy tasks with high data efficiency. DeepSeek-R1 [14] is an enhanced reasoning model designed to improve LLMs’ reasoning performance that incorporates multi-stage training and cold-start data before the large-scale RL. DeepSeek-R1 demonstrates remarkable reasoning capabilities, and achieves comparable performance to OpenAI-o1 across various reasoning tasks, such as mathematical problems, code generation, and scientific reasoning. Additionally, there are also open-sourced medium-sized LLMs with strong reasoning capabilities, e.g., DeepSeek-R1-32B, a distilled version of DeepSeek-R1, QwQ-32B [27], Gemma 3 [26].

### v2

@@ Abstract
Reasoning is the fundamental capability of large language models (LLMs). Due to the rapid progress of LLMs, there are two main issues of current benchmarks: i) these benchmarks can be crushed in a short time (less than 1 year), and ii) these benchmarks may be easily hacked. To handle these issues, we propose the ever-scalingness for building the benchmarks which are scaling over complexity against crushing, instance against hacking and exploitation, oversight for easy verification, and coverage for real-world relevance. This paper presents Nondeterministic Polynomial-time Problem Challenge (NPPC), an ever-scaling reasoning benchmark for LLMs. Specifically, the NPPC has three main modules: i) npgym, which provides a unified interface of 25 well-known NP-complete problems and can generate any number of instances with any levels of complexities, ii) npsolver, which provides a unified interface to evaluate the problem instances with both online and offline models via APIs and local deployments, respectively, and iii) npeval, which provides the comprehensive and ready-to-use tools to analyze the performances of LLMs over different problems, the number of tokens, the reasoning errors and the solution errors. Extensive experiments over widely-used LLMs demonstrate: i) NPPC can successfully decrease the performances of advanced LLMs to below 10%, demonstrating that NPPC is not crushed by current models, ii) DeepSeek-R1, Claude-3.7-Sonnet, and o1/o3-mini are the most powerful LLMs, where DeepSeek-R1 can outperform Claude-3.7-Sonnet and o1/o3-mini in most NP-complete problems considered, and iii) the numbers of tokens in the advanced LLMs, e.g., Claude-3.7-Sonnet and DeepSeek-R1, are observed first to increase and then decrease when the problem instances become more and more difficult. Through continuously scaling analysis, NPPC can provide critical insights into the limits of LLMs’ reasoning capabilities, exposing fundamental limitations and suggesting future directions for further improvements.
@@ 1 Introduction
The remarkable successes of Large Language Models (LLMs) (Achiam et al., 2023) have catalyzed the fundamental shift of artificial intelligence. Recent breakthroughs on reasoning (Guo et al., 2025) enable LLMs to complete complex tasks, e.g., math proof, code generation and computer use, which require the capabilities of understanding, generation and long-term planning. Various benchmarks, e.g., GPQA (Rein et al., 2024), AIME, SWE-bench (Jimenez et al., 2024) and ARC-AGI (Chollet, 2019), are proposed to evaluate these advanced reasoning capabilities, where most benchmarks are curated and verified by human researchers with a finite number of questions. These benchmarks guide the directions for advancing LLM capabilities.
@@
Current benchmarks face two fundamental challenges that limit their effectiveness for LLM evaluation. First, current benchmarks can be crushed in a short time: GSM8K (Cobbe et al., 2021) performance increased from approximately 35% to 95% within three years, while SWE-bench (Jimenez et al., 2024) scores improved from 7.0% to 64.6% in merely eight months, as illustrated in Figure 1. This rapid saturation suggests that these benchmarks quickly lose their discriminative power as models advance. Second, current benchmarks can be easily hacked or exploited. Static benchmarks are susceptible to data contamination and memorization issues, leading to overfitting rather than genuine capability assessment (Wu et al., 2025; Xu et al., 2024). While live benchmarks such as LiveCodeBench (Jain et al., 2025) address contamination by continuously introducing new problems, they require substantial ongoing human curation efforts. Similarly, human evaluation platforms like ChatbotArena (Chiang et al., 2024) incur significant costs (approximately \$3,000 per evaluation) and remain vulnerable to strategic manipulation where MixEval (Ni et al., 2024) can achieve comparable correlation with human judgment at under \$1 per evaluation. These limitations represent significant obstacles for reliable evaluation of the rapidly evolved LLMs.
Figure 2: Desiderate of ever-scalingness
To address these issues, we propose the ever-scalingness with four desiderata for a benchmark (as shown in Figure 2): i) scaling over complexity – the benchmark can generate the problems with continually increasing complexities to avoid the crushing of the benchmarks. This ensures that as LLMs improve, the benchmark remains challenging by providing harder problems that push the boundaries of current capabilities for long-term differentiation. ii) scaling over instance – the benchmark can generate an infinite number of instances to avoid the exploitation and overfitting. This prevents LLMs from memorizing specific problem instances during training and ensures that the evaluation truly measures the reasoning capabilities of LLMs. iii) scaling over oversight – the benchmark can verify the correctness of the solutions efficiently for the problems with any complexity. This is crucial because as problem complexity increases, manual verification becomes impractical or impossible. Automated and reliable verification enables evaluation at scale and ensures that harder problems can be verified efficiently, making the benchmark practically sustainable. iv) scaling over coverage – the benchmark should comprehensively cover problem types that are highly relevant to real-world applications, rather than focusing on puzzles. or rare edge cases. This ensures that performance improvements on the benchmark translate to practical value, measuring capabilities that matter for downstream tasks and real-world deployment rather than narrow, specialized skills. These four desiderata for ever-scalingness ensure continuous differentiation among LLMs over extended periods, identifying fundamental limitations for further improvement.
To construct the ever-scaling benchmark, we focus on nondeterministic polynomial-time (NP) problems whose solutions can be verified in polynomial time (Cormen et al., 2022). Specifically, we target on NP-complete (NPC) problems, i.e., the most computationally challenging problems in the NP class, for three key reasons. First, NPC problem instances can be systematically generated across arbitrary difficulty levels through controlled parameters, e.g., numbers of variables, enabling precise scaling of both complexity and instance. Second, NPC problems are intrinsically “difficult to solve, easy to verify”, i.e., no polynomial-time algorithms have been discovered for solving NPC problems, making them computationally intractable even with specialized tools, while their solutions remain efficiently verifiable. Problems in the P complexity class can be solved in polynomial time. When LLMs are equipped with code execution capabilities, they can generate and execute algorithms to solve these problems directly. Therefore, such benchmarks become susceptible to trivial solutions through computational tools rather than genuine reasoning. Conversely, NP-hard problems, particularly those lacking polynomial-time verification procedures, present the challenge for verifying the solutions for large-scale problem instances. Third, NPC problems demonstrate broad applicability, including diverse real-world scenarios (e.g., routing (Toth and Vigo, 2002) and protein folding (Crescenzi et al., 1998)) and various puzzles, e.g., Sudoku (Seely et al., 2025). The theoretical foundation for using NPC problems as a comprehensive evaluation framework stems from the fundamental property that any NP problem can be reduced to an NPC problem in polynomial time, establishing NPC problems as a theoretically grounded, universal framework for computational problem-solving assessment. Therefore, NPC problems are the foundation problems of all computational problems and LLMs are the foundation models for wide range tasks, thus leading to our ever-scaling nondeterministic polynomial-time problem challenge (NPPC) (Figure 4(a)).
@@
Specifically, NPPC has three main modules: i) npgym, which provides a unified interface of 25 well-known NPC problems and can generate any number of instances with any levels of complexities, which implies the ever-scalingness of NPPC, ii) npsolver, which provides a unified interface to evaluate the problem instances with both online and offline models via APIs and local deployments, respectively, to facilitate users to evaluate their own models and iii) npeval, which provides comprehensive and ready-to-use tools to analyze the performances of LLMs over different problems, the number of tokens, the “aha moments”, the reasoning errors and the solution errors, which can provide in-depth analysis of the LLMs and the insights to further improve the LLMs’ reasoning capabilities. Extensive experiments over widely-used LLMs, i.e., GPT-4o-mini, GPT-4o, Claude-3.7-Sonnet, DeepSeek-V3, DeepSeek-R1, and OpenAI o1-mini, demonstrate: i) NPPC can successfully decrease the performances of advanced LLMs to below 10%, demonstrating that NPPC is not crushed by current LLMs, ii) DeepSeek-R1, Claude-3.7-Sonnet, and o1/o3-mini are the most powerful LLMs, where DeepSeek-R1 can outperform Claude-3.7-Sonnet and o1-mini in most NP-complete problems considered, and iii) the numbers of tokens in the advanced LLMs, e.g.. Claude-3.7-Sonnet and DeepSeek-R1, are observed to first increase and then decrease when the problem instances become more and more difficult. We also analyze the typical reasoning errors in the LLMs, which provide the insights of the fundamental limitations of current LLMs and suggest the potential directions for further improvement. To the best of our knowledge, NPPC is the first ever-scaling benchmark for reliable and rigorous evaluation of the reasoning limits of LLMs, according to the four desiderata defined previously.
@@ 2 Related Work
Complexity
Instance
Oversight
Coverage
NPHardEval (Fan et al., 2024)
ZebraLogic (Lin et al., 2025)
Reasoning Gym (Stojanovski et al., 2025)
Sudoku-Bench (Seely et al., 2025)
ARC-AGI-1 & 2 (Chollet, 2019)
@@
Traditional benchmarks are typically curated by human with static datasets. Abstraction and Reasoning Corpus (ARC-AGI)-1 (Chollet, 2019) is designed to be “easy for humans, hard for AI”, which is formed by human-curated 800 puzzle-like tasks, designed as grid-based visual reasoning problems. o3 at high compute scored 87% on ARC-AGI-1 (OpenAI, 2025), which roughly crushes the ARC-AGI-1 benchmarks and leads to the emergence of the ARC-AGI-2 benchmark. This pattern exemplifies a fundamental challenge with traditional benchmarks for LLMs, including MMLU (Hendrycks et al., 2021), GPQA (Rein et al., 2024), GSM8K (Cobbe et al., 2021), and SWE-bench (Jimenez et al., 2024), where static benchmarks are systematically solved within relatively short periods (as shown in Figure 1). Therefore, researchers have to continuously either develop new benchmarks, e.g., MMLU-Pro (Wang et al., 2024b) and SuperGPQA (Du et al., 2025), or regularly update with new datasets and problems, e.g., LiveCodeBench (Jain et al., 2025) and SWE-bench-Live (Zhang et al., 2025). However, these remedies rely on extensive human efforts to maintain their relevance and difficulty and cannot fully address the crushing issue of benchmarks.
Several recent benchmarks consider either NP(C) problems, e.g., 3SAT (Balachandran et al., 2025; Hazra et al., 2024; Parashar et al., 2025), or partially the ever-scalingness (Fan et al., 2024; Stojanovski et al., 2025) (displayed in Table 1). NPHardEval (Fan et al., 2024) considers 3 problems from P, NPC and NP-hard classes and use these class to evaluate the LLMs. We note that the problems in P class can be solved by augmenting the LLMs with tools, e.g., code running, and the NP-hard problems cannot be verified efficiently, therefore, NPHardEval cannot scale over the scalable oversight. Only 3 NPC problems are considered, i.e., Knapsack, Traveling salesman problem (TSP) and graph coloring, and the instances of each problem in NPHardEval are finite and only regularly updated, which cannot scale over the instance and complexity. ZebraLogic (Lin et al., 2025) considers one logic puzzle, i.e., Zebra puzzle, to test the reasoning capabilities of LMs when the problems’ complexities increase. However, the reasoning capability on specific puzzles does not necessarily transfer to other problems, which violates the scaling of the coverage. Sudoku-Bench (Seely et al., 2025) focuses on one specific Sudoku game with 2765 procedurally generated instances with various difficulty levels. Reasoning Gym (Stojanovski et al., 2025) is an ongoing project which collects the procedural generators and algorithmic verifiers for infinite training data with adjustable complexity. Though with some NP(C) problems, e.g., Zebra puzzles and Sudoku, the reasoning gym does not specifically focus on NPC problems and cannot meet the desiderata of ever-scalingness. Several recent work leverages NP(C) problems as the training environments to improve the general reasoning capabilities of LLMs (Zeng et al., 2025; Li et al., 2025; Liu et al., 2025a), which demonstrates the advantages of NP(C) problems in terms of controllable difficulty and efficient solution verification. These existing efforts highlight the need for a systematic approach for training and benchmarking LLMs grounded in the complexity theory. This motivates the formulation of ever-scalingness and the development of NPPC.
@@ 3 Preliminaries
P and NP Problems. The problems in P class are decision problems that can be solved in polynomial time by a deterministic Turing machine, which implies there exists an algorithm that can find a solution in time proportional to a polynomial function, e.g., $O(n^{k})$, of the input size $n$. Examples include sorting, shortest path problems, and determining if a number is prime. The problems in NP class are decision problems that can be solved in polynomial time by nondeterministic Turing machine, where a proposed solution can be easily verified, though finding that solution might require more time (as displayed in Definition 1). All P problems are also in NP, but the reverse remains an open question, known as “P vs. NP problem”. NP problems form the cornerstone of computational complexity theory, for which solution verification is tractable (polynomial time) even though solution discovery may be intractable (potentially exponential time), i.e., “difficult to solve, easy to verify”. Many real-world optimization problems can be formulated as NP problems, such as equilibrium finding in game theory, portfolio management, network design and machine learning.
@@
NP-complete (NPC) Problems. Formally, a problem $\Omega$ is an NPC problem if i) the problem is in NP, and ii) any NP problems can be transformed to problem $\Omega$ in polynomial time. This reducibility property establishes NPC problems as the “hardest” problems in NP class. The Cook-Levin theorem established SAT as the first proven NPC problem (Cook, 2023; Karp, 2009), while 3SAT is the special case of SAT and is also an NPC problem. Subsequent NPC problems typically proven via reduction chains back to 3SAT or other established NPC problems. The most well-known NPC problems include vertex cover problem, clique problem, traveling salesman proble (TSP), Hamiltonian path/cycle problem, etc. NPC problems play the most important roles in answering the “P vs. NP problem”, i.e., if any NPC problem were shown to have a polynomial-time algorithm, then P = NP. However, despite decades of research, no polynomial-time algorithms for any NPC problem is discovered, which implies that NPC problems are computationally intractable by current methods. While NP-hard problems represent a broader class that includes optimization variants and potentially harder problems, they are less suitable for benchmarks because their solutions cannot be verified in polynomial time, which fundamentally limits the scaling of complexity and oversight of the benchmarks.
Reasoning in LLMs. The reasoning ability of LLMs refers to the model’s capacity to tackle complex problems, e.g., mathematical proof, code generation through multi-step thinking and context understanding. Recently, specialized reasoning models have been proposed. OpenAI-o1 is an LLM trained with reinforcement learning (RL), which enables the model to perform complex reasoning, including logical thinking and problem solving, via chain-of-thought (CoT). o1 thinks before it answers and can significantly outperform GPT‑4o on reasoning-heavy tasks with high data efficiency. DeepSeek-R1 (Guo et al., 2025) is an enhanced reasoning model designed to improve LLMs’ reasoning performance that incorporates multi-stage training and cold-start data before the large-scale RL. DeepSeek-R1 demonstrates remarkable reasoning capabilities, and achieves comparable performance to OpenAI-o1 across various reasoning tasks, e.g., mathematical problems, code generation, and scientific reasoning. Additionally, there are open-sourced medium-sized LLMs with strong reasoning capabilities, e.g., DeepSeek-R1-32B, a distilled version of DeepSeek-R1, and QwQ-32B.